# SLRQA: A Sparse Low-Rank Quaternion Model for Color Image Processing with Convergence Analysis

### v1

@@ 4 Convergence analysis
$\nabla g(\mathbf{\dot{W}}_{k})+\beta\mathcal{B}^{\#}(\mathcal{C}{\mathbf{\dot{X}}}_{k+1}-\mathcal{B}\mathbf{\dot{W}}_{k+1}-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta)+L_{2,k}\Delta\mathbf{\dot{W}}_{k+1}=0.$ (40)
@@
$\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k+1}}$ $=\nabla g(\mathbf{\dot{W}}_{k+1})-\nabla g(\mathbf{\dot{W}}_{k})+\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{2,k}\Delta\mathbf{\dot{W}}_{k+1}\in\partial_{\mathbf{\dot{W}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1}).$ (41)
@@
Since $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ is bounded by Theorem 4.1. There exists at least one limit point of $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ and $\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)$ is bounded. According to [27, Chapter 2, Exercise 6], it follows that $\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)$ is closed. Hence $\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)$ is nonempty and compact. As a consequence, $\lim_{k\rightarrow\infty}\text{dist}\left[({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k}),\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)\right]=0.$ It follows from Lemma 4.7 that $\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)\subseteq\text{crit}\,\mathcal{L}_{\beta}$, which completes the proof.
@@
$\mathbf{\dot{S}}_{{\mathbf{\dot{X}}}_{k}}:=\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k}},\quad\mathbf{\dot{S}}_{\mathbf{\dot{W}}_{k}}:=\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}}+2r\theta_{0}\Delta\mathbf{\dot{W}}_{k},\quad\mathbf{\dot{S}}_{\mathbf{\dot{\Lambda}}_{k}}:=\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k}}+2r\theta_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k},$
$\mathbf{\dot{S}}_{{\mathbf{\dot{W}}^{\prime}}_{k}}:=-2r\theta_{0}\Delta\mathbf{\dot{W}}_{k},\quad\mathbf{\dot{S}}_{{\mathbf{\dot{\Lambda}}^{\prime}}_{k}}:=-2\theta_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}$
@@
$\tilde{\pi}=\sqrt{3}\pi+4r\max\{\theta_{0},\theta_{2}\|\mathcal{B}\|^{2}\},$ (44)
@@
$\mathbf{\dot{S}}_{\mathbf{\dot{W}}_{k}}$ ::=:= $\nabla_{\mathbf{\dot{W}}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=\nabla_{\mathbf{\dot{W}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})+2r\theta_{0}\Delta\mathbf{\dot{W}}_{k}=\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}}+2r\theta_{0}\Delta\mathbf{\dot{W}}_{k},$
$\mathbf{\dot{S}}_{\mathbf{\dot{\Lambda}}_{k}}$ ::=:= $\nabla_{\mathbf{\dot{\Lambda}}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=\nabla_{\mathbf{\dot{\Lambda}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})+2r\theta_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}=\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k}}+2r\theta_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k},$
$\mathbf{\dot{S}}_{{\mathbf{\dot{W}}^{\prime}}_{k}}$ ::=:= $\nabla_{\mathbf{\dot{W}}^{\prime}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=-2r\theta_{0}\Delta\mathbf{\dot{W}}_{k},$
$\mathbf{\dot{S}}_{{\mathbf{\dot{\Lambda}}^{\prime}}_{k}}$ ::=:= $\nabla_{\mathbf{\dot{\Lambda}}^{\prime}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=-2r\theta_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}.$
@@
$\|\mathbf{\dot{S}}_{{\mathbf{\dot{X}}}_{k}}\|_{\mathrm{F}}=\|\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k}}\|_{\mathrm{F}},\quad\|\mathbf{\dot{S}}_{\mathbf{\dot{W}}_{k}}\|_{\mathrm{F}}\leq\|\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}}\|_{\mathrm{F}}+2r\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}},$
$\|\mathbf{\dot{S}}_{\mathbf{\dot{\Lambda}}_{k}}\|_{\mathrm{F}}\leq\|\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k}}\|_{\mathrm{F}}+2r\theta_{2}\|\mathcal{B}\|^{2}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}},\quad\|\mathbf{\dot{S}}_{{\mathbf{\dot{W}}^{\prime}}_{k}}\|_{\mathrm{F}}=2r\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}},\quad\|\mathbf{\dot{S}}_{{\mathbf{\dot{\Lambda}}^{\prime}}_{k}}\|_{\mathrm{F}}=2r\theta_{2}\|\mathcal{B}\|^{2}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}.$
@@
$\leq$ $\|\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k}}\|_{\mathrm{F}}+\|\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}}\|_{\mathrm{F}}+\|\mathbf{\dot{D}}^{k}_{\mathbf{\dot{\Lambda}}}\|_{\mathrm{F}}+4r\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+4r\theta_{2}\|\mathcal{B}\|^{2}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}$
$\leq$ $\sqrt{3}|||\mathbf{\dot{D}}_{k}|||+4r\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+4r\theta_{2}\|\mathcal{B}\|^{2}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}$
$\leq$ $\sqrt{3}\rho\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}+(\sqrt{3}\rho+4r\theta_{0})\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+(\sqrt{3}\rho+4r\theta_{2}\|\mathcal{B}\|^{2})\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}},$
@@
The following convergence result of the whole sequence, established under the Kurdyka-Łojasiewicz property, follows from [36, Theorem 2]. The difference lies in the construction of $\mathcal{R}_{k}$. The proofs are given for completeness.
@@
Suppose the penalty parameter $\beta$ is large enough, and $L_{1,k},L_{2,k}$ are set as given constants for all $k$ such that (20) and Assumption (A5) hold, then the sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ generated by Algorithm 1 converges to a stationary point of SLRQA (2).
@@
It follows directly from the definition of SLRQA (2) that Assumption A1,A2 and A3 are all satisfied. Since (20) and Assumption (A5) hold, Assumption A4, A5 are also satisfied. Furthermore, the Huber function and Frobenius norm are semi-algebraic [3, Example 2]. From [28, Appendix, Lemma 4], it follows that the nuclear norm is also a semi-algebraic function. Since $\phi$ is a semi-algebraic function and the fact that the composition and sum of semi-algebraic functions are also semi-algebraic, we can conclude that $\mathcal{R}$ is semi-algebraic and the KL-inequality 46 holds since any proper closed semi-algebraic function satisfies the KL-inequality [3, Theorem 3]. It follows from Theorem 4.4 that the sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ generated by Algorithm 1 converges to a stationary point of SLRQA (2), which completes the proof. ∎
5 An Algorithm for SLRQA-NF model
In this section, we propose the PL-ADMM-NF algorithm to solve SLRQA-NF (4) which is summarized in Algorithm 3. It is noted that SLRQA-NF is a model of the limiting case of SLRQA. The linearization technique is applied to the update of $\mathbf{\dot{W}}$. Each iteration of PL-ADMM-NF consists of minimizing the augmented Lagrange function with respect to $\mathbf{\dot{W}}$ variable after linearization, minimizing the augmented Lagrange function with respect to ${\mathbf{\dot{X}}}$ variable, and updating the multiplier.
Algorithm 3 PL-ADMM-NF algorithm for SLRQA-NF in (4)
1:Corrupted images $\mathbf{\dot{Y}}$; two parameters $\mu\in(0,2)$ and $\beta>0$; a tolerance $\eta>0$; a sequence of regularization parameters $\{L_{1,k}\}$ satisfying $\sup_{k\geq 0}{L_{1,k}}<\infty$ and $\inf_{k>0}{L_{1,k}}>0$.
2:Recovered $\mathbf{\dot{Z}}_{*}$.
4:for $k=0,1,2,\ldots$ do
5:     $\mathbf{\dot{Z}}$ subproblem:
$\mathbf{\dot{Z}}_{k+1}\in\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{Z}}}\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Z}}),\gamma)+\left\langle\mathbf{\dot{\Lambda}}_{1,k},\mathbf{\dot{Z}}-\mathbf{\dot{W}}_{k}\right\rangle+\frac{\beta}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}+\frac{L_{1,k}}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{Z}}_{k}\|_{\mathrm{F}}^{2};$
6:     $\mathbf{\dot{W}}$ subproblem:
$\mathbf{\dot{W}}_{k+1}$ $=\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{W}}}\lambda\left\langle\nabla p(\mathbf{\dot{W}}_{k}),\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\rangle+\beta\left\langle\mathbf{\dot{W}}_{k}-\mathbf{\dot{Z}}_{k+1}+\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{1,k},\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\rangle$
$+\beta\left\langle\mathcal{W}\mathcal{A}^{\#}(\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{Y}})+\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{2,k},\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\rangle+\frac{\beta\eta}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2};$
7:     $\mathbf{\dot{\Lambda}}_{1,k+1}=\mathbf{\dot{\Lambda}}_{1,k}+\mu\beta\cdot(\mathbf{\dot{Z}}_{k+1}-\mathbf{\dot{W}}_{k+1})$;
8:     $\mathbf{\dot{\Lambda}}_{2,k+1}=\mathbf{\dot{\Lambda}}_{2,k}+\mu\beta\cdot\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{Y}})$;
10:         $\mathbf{\dot{Z}}_{*}=\mathbf{\dot{Z}}_{k+1}$; break
11:     end if
12:end for

### v2

@@ 4.1 Global Convergence for Algorithm 1
$\nabla g(\mathbf{\dot{W}}_{k})+\beta\mathcal{B}^{\#}(\mathcal{C}{\mathbf{\dot{X}}}_{k+1}-\mathcal{B}\mathbf{\dot{W}}_{k+1}-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta)+L_{k,2}\Delta\mathbf{\dot{W}}_{k+1}=0.$ (46)
@@
$\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k+1}}$ $=\nabla g(\mathbf{\dot{W}}_{k+1})-\nabla g(\mathbf{\dot{W}}_{k})+\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{k,2}\Delta\mathbf{\dot{W}}_{k+1}\in\partial_{\mathbf{\dot{W}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1}).$ (47)
@@
Since $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ is bounded by Theorem 4.1. There exists at least one limit point of $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ and $\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)$ is bounded. According to [32, Chapter 2, Exercise 6], it follows that $\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)$ is closed. Hence $\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)$ is nonempty and compact. As a consequence, $\lim_{k\rightarrow\infty}\text{dist}\left[({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k}),\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)\right]=0.$ It follows from Lemma 4.7 that $\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)\subseteq\text{crit}\,\mathcal{L}_{\beta}$, which completes the proof.
@@
$\mathbf{\dot{S}}_{{\mathbf{\dot{X}}}_{k}}:=\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k}},\quad\mathbf{\dot{S}}_{\mathbf{\dot{W}}_{k}}:=\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}}+2r\varsigma_{0}\Delta\mathbf{\dot{W}}_{k},\quad\mathbf{\dot{S}}_{\mathbf{\dot{\Lambda}}_{k}}:=\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k}}+2r\varsigma_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k},$
$\mathbf{\dot{S}}_{{\mathbf{\dot{W}}^{\prime}}_{k}}:=-2r\varsigma_{0}\Delta\mathbf{\dot{W}}_{k},\quad\mathbf{\dot{S}}_{{\mathbf{\dot{\Lambda}}^{\prime}}_{k}}:=-2\varsigma_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}$
@@
$\tilde{\pi}=\sqrt{3}\pi+4r\max\{\varsigma_{0},\varsigma_{2}\|\mathcal{B}\|^{2}\},$ (50)
@@
$\mathbf{\dot{S}}_{\mathbf{\dot{W}}_{k}}$ ::=:= $\nabla_{\mathbf{\dot{W}}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=\nabla_{\mathbf{\dot{W}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})+2r\varsigma_{0}\Delta\mathbf{\dot{W}}_{k}=\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}}+2r\varsigma_{0}\Delta\mathbf{\dot{W}}_{k},$
$\mathbf{\dot{S}}_{\mathbf{\dot{\Lambda}}_{k}}$ ::=:= $\nabla_{\mathbf{\dot{\Lambda}}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=\nabla_{\mathbf{\dot{\Lambda}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})+2r\varsigma_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}=\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k}}+2r\varsigma_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k},$
$\mathbf{\dot{S}}_{{\mathbf{\dot{W}}^{\prime}}_{k}}$ ::=:= $\nabla_{\mathbf{\dot{W}}^{\prime}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=-2r\varsigma_{0}\Delta\mathbf{\dot{W}}_{k},$
$\mathbf{\dot{S}}_{{\mathbf{\dot{\Lambda}}^{\prime}}_{k}}$ ::=:= $\nabla_{\mathbf{\dot{\Lambda}}^{\prime}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=-2r\varsigma_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}.$
@@
$\|\mathbf{\dot{S}}_{{\mathbf{\dot{X}}}_{k}}\|_{\mathrm{F}}=\|\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k}}\|_{\mathrm{F}},\quad\|\mathbf{\dot{S}}_{\mathbf{\dot{W}}_{k}}\|_{\mathrm{F}}\leq\|\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}}\|_{\mathrm{F}}+2r\varsigma_{0}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}},$
$\|\mathbf{\dot{S}}_{\mathbf{\dot{\Lambda}}_{k}}\|_{\mathrm{F}}\leq\|\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k}}\|_{\mathrm{F}}+2r\varsigma_{2}\|\mathcal{B}\|^{2}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}},\quad\|\mathbf{\dot{S}}_{{\mathbf{\dot{W}}^{\prime}}_{k}}\|_{\mathrm{F}}=2r\varsigma_{0}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}},\quad\|\mathbf{\dot{S}}_{{\mathbf{\dot{\Lambda}}^{\prime}}_{k}}\|_{\mathrm{F}}=2r\varsigma_{2}\|\mathcal{B}\|^{2}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}.$
@@
$\leq$ $\|\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k}}\|_{\mathrm{F}}+\|\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}}\|_{\mathrm{F}}+\|\mathbf{\dot{D}}^{k}_{\mathbf{\dot{\Lambda}}}\|_{\mathrm{F}}+4r\varsigma_{0}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+4r\varsigma_{2}\|\mathcal{B}\|^{2}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}$
$\leq$ $\sqrt{3}|||\mathbf{\dot{D}}_{k}|||+4r\varsigma_{0}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+4r\varsigma_{2}\|\mathcal{B}\|^{2}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}$
$\leq$ $\sqrt{3}\rho\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}+(\sqrt{3}\rho+4r\varsigma_{0})\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+(\sqrt{3}\rho+4r\varsigma_{2}\|\mathcal{B}\|^{2})\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}},$
@@
The following convergence result of the whole sequence, established under the Kurdyka-Łojasiewicz(KL) property, follows from [43, Theorem 2]. The difference lies in the construction of $\mathcal{R}_{k}$. The proofs are given for completeness. Furthermore, the rationality of the KL property (52) assumption for Model (2) is clarified in Corollary 4.1.
@@
Suppose the penalty parameter $\beta$ is large enough, and $L_{k,1},L_{k,2}$ are set as given constants for all $k$ such that (26) and Assumption (A5) hold, then the sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ generated by Algorithm 1 converges to a stationary point of SLRQA (2).
@@
It follows directly from the definition of SLRQA (2) that Assumption A1,A2 and A3 are all satisfied. Since (26) and Assumption (A5) hold, Assumption A4, A5 are also satisfied. Furthermore, the Huber function and Frobenius norm are semi-algebraic [5, Example 2]. From [34, Appendix, Lemma 4], it follows that the nuclear norm is also a semi-algebraic function. Since $\phi$ is a semi-algebraic function and the fact that the composition and sum of semi-algebraic functions are also semi-algebraic, we can conclude that $\mathcal{R}$ is semi-algebraic and the KL-inequality 52 holds since any proper closed semi-algebraic function satisfies the KL-inequality [5, Theorem 3]. It follows from Theorem 4.4 that the sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ generated by Algorithm 1 converges to a stationary point of SLRQA (2), which completes the proof. ∎
4.2 Global Convergence for Algorithm 3